OpenStudy (anonymous):

Solve 6^5x = 20 Round to the nearest ten-thousandth x=_______

OpenStudy (anonymous):

satellite you are a bro...

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] by the change of base formula, which is what you need here

OpenStudy (anonymous):

can you plug it in for me?

OpenStudy (anonymous):

you get in one step \[5x=\frac{\ln(20)}{\ln(6)}\] making \[x=\frac{\ln(20)}{5\ln(6)}\] and a calculator